OpenStudy (anonymous):

Find the radius of convergence of the series:

OpenStudy (anonymous):

\[\large \sum_{n=1}^{\infty} \frac{n!x^n}{n^n}\] What I did was apply the root test, so: \[\huge \lim_{n \rightarrow \infty}\sqrt[n]{\frac{n!x^n}{n^n}}= x \lim_{n \rightarrow \infty} \frac{\sqrt[n]{n!}}{n}\] I applied L'Hospital's rule and determined that the limit is 0, which means the radius should be \(\infty\), but that's not the right answer =/

OpenStudy (shadowlegendx):

Iʻd ask @Hero

OpenStudy (anonymous):

Oh wait, that just means that it's convergent, right? Since L<0. It's been a while since I've taken Calculus 2 >_< Taking differential equations right now..

OpenStudy (anonymous):

L<1*

OpenStudy (thomas5267):

http://www.wolframalpha.com/input/?i=Lim [%28%282+Pi+n%29^%281%2F%282n%29%29%29%2Fe%2Cn-%3EInfinity] \[ \lim_{x\to\infty}\frac{2\pi n^{\frac{1}{2n}}}{e}=\frac{1}{e} \] Don't ask me why.

OpenStudy (thomas5267):

Put this into WolframAlpha. Lim[((2 Pi n)^(1/(2n)))/e,n->Infinity]

OpenStudy (anonymous):

That doesn't make sense to me though. I don't necessarily want the answer -- I want to know how to solve it.

OpenStudy (zarkon):

use ratio test

OpenStudy (anonymous):

@Zarkon Shouldn't they give me the same answer?

OpenStudy (zarkon):

yes...ratio test is the method to use for this problem though

OpenStudy (zarkon):

otherwise you need to use Stirling's approximation

OpenStudy (thomas5267):

\[ \begin{align*} &\phantom{{}={}}\lim_{n\to\infty}\sqrt[n]{\frac{n!}{n^n}}\\ &=\lim_{n\to\infty}\sqrt[n]{\frac{\sqrt{2\pi n}(\frac{n}{e})^n}{n^n}}\\ &=\lim_{n\to\infty}\frac{(2\pi n)^\frac{1}{2n}(\frac{n}{e})}{n}\\ &=\lim_{n\to\infty}\frac{(2\pi n)^\frac{1}{2n}}{e}\\ &=\frac{1}{e}\lim_{n\to\infty}n^\frac{1}{2n} \end{align*} \] Take logarithm of that limit than solve it.

OpenStudy (anonymous):

I still got 0 =/ `Ratio Test`: \[\large \lim_{n \rightarrow \infty}\frac{ (n+1)!x^{n+1} }{ (n+1)^{n+1} }* \frac{ n^n }{ n!x^n }=\lim_{n \rightarrow \infty}\frac{ n }{ (n+1)^n }\]\[\large L'Hospital's \rightarrow \lim_{n \rightarrow \infty}\frac{ 1 }{ n(n+1)^{n-1} }=0\]

OpenStudy (zarkon):

that is because you did it wrong

OpenStudy (anonymous):

Oh, where did I mess up?

OpenStudy (zarkon):

\[\frac{ (n+1)!x^{n+1} }{ (n+1)^{n+1} }\frac{ n^n }{ n!x^n }=\frac{(n+1)xn^n}{(n+1)^{n+1}}=\frac{xn^n}{(n+1)^{n}}\]

OpenStudy (anonymous):

Ah yeah I left out the exponent of n on the top.. But how would I take the limit of that then?

OpenStudy (zarkon):

\[\frac{xn^n}{(n+1)^{n}}=\frac{x}{\frac{(n+1)^n}{n^n}}=\frac{x}{\left(\frac{n+1}{n}\right)^n}=\frac{x}{\left(1+\frac{1}{n}\right)^n}\]

OpenStudy (zarkon):

\[\lim_{n\to\infty}\left(1+\frac{1}{n}\right)^n=e\]

OpenStudy (anonymous):

Why does that equal e? Wouldn't it be 1 then? 1/infinity (approaching it) would be 0. so (1+0)^infinity would be 1?

OpenStudy (zarkon):

http://mathworld.wolfram.com/e.html see equation (3). it is a definition of e

OpenStudy (anonymous):

Ah okay, I was never taught that. Definitely good to know now though. Thanks!!